### v3

# InteLiPlan: An Interactive Lightweight LLM-Based Planner for Domestic Robot Autonomy

Abstract
We introduce an interactive LLM-based framework designed to enhance the autonomy and robustness of domestic robots, targeting embodied intelligence. Our approach reduces reliance on large-scale data and incorporates a robot-agnostic pipeline that embodies an LLM. Our framework, InteLiPlan, ensures that the LLM’s decision-making capabilities are effectively aligned with robotic functions, enhancing operational robustness and adaptability, while our human-in-the-loop mechanism allows for real-time human intervention when user instruction is required. We evaluate our method in both simulation and on the real Toyota Human Support Robot and Anymal D-Unitree Z1 platforms. Our method achieves a 95% success rate in the ‘fetch me’ task completion with failure recovery, highlighting its capability in both failure reasoning and task planning. InteLiPlan achieves comparable performance to state-of-the-art large-scale LLM-based robotics planners, while using only real-time onboard computing.
I Introduction
In recent years, incorporating large language models (LLMs) and vision-language models (VLMs) into robotic systems as high-level planners has made remarkable progress in the embodied AI field, with applications in domestic autonomy and robotic assistance. In such hierarchical frameworks, multimodal perception modules are integrated into the system, and LLMs or VLMs serve as central components for scene understanding and reasoning, as well as generating action steps. Exploiting the reasoning abilities of these advanced models, LLM-based frameworks have shown promise in enhancing human-robot interaction and enabling intuitive decision-making in robotics.
Despite their impressive capabilities, applying LLMs in robotics presents unique challenges [1]. These challenges stem primarily from the constraints of robot kinematics and the dynamic nature of the environments in which robots operate. Additionally, integrating these generative models within the existing robotics pipelines is non-trivial, as these are usually tightly coupled systems of perception, planning, and actuation. Having such systems running in real-time robotic deployment also poses a significant challenge due to the computational overhead. Moreover, robotic motion safety and robustness are critical concerns in human-robot environments. Therefore, developing a general-purpose autonomous robotic system for domestic scenarios remains a challenging and open problem.
Figure 1: Step-by-step execution of InteLiPlan result on the physical Anymal D-Unitree Z1. In our system, the robot receives either requests or guidance from humans. Our lightweight onboard planner will generate the robot actions and execute them sequentially.
In this work, we propose a lightweight LLM-based framework designed for domestic robots, aiming to address the aforementioned challenges. Our approach reduces reliance on large-scale data and develops a robot-agnostic pipeline. Furthermore, we detail strategies to mitigate kinematic constraints through real-time feasibility checks, ensuring that decision-making processes facilitated by the LLM align effectively with the physical capabilities of robots.
An integral part of our framework is the human-in-the-loop (HITL) mechanism, which allows the system to handle system failures (e.g., partially observable environment, vision failures due to lightning, etc.) or task confusions through human intervention. This feature is critical for maintaining autonomy, especially in complex environments where decisions by the LLM may fail to reason what the user requested from an embodied perspective. The system allows human operators to input corrective actions directly, facilitating on-the-fly adaptation to unanticipated situations. This real-time interaction not only enhances the robustness and reliability of robotic operations but also enables adjusting the LLM’s responses in accordance with human guidance.
By tackling these critical issues, our framework not only enhances the practical applicability of LLMs in robotics but also paves the way for more intuitive and responsive robotic systems capable of complex interactions and tasks in varied real-world scenarios. In summary, our work investigates the following questions:
  • How can we design a lightweight LLM-based planner that effectively integrates user commands, visual perception, and action feasibility with onboard computing resources?
  • How can robots make use of LLMs to translate human-language inputs to reliable and executable action directly?
  • How can uncertainties in robotic planning be solved through human intervention?
  • How can reachability-aware LLM-based robotic planners be applicable cross-embodiment?
II Related work
II-A LLMs for Robotic Systems
The application of LLM models in robotics, driven by recent advancements in natural language processing, has become a focal point of current research [2]. Silver et al. [3] study how a pretrained LLM can solve planning domain definition language (PDDL) problems. Similarly, Plansformer [4] requires PDDL domain inputs and fine-tunes the transformer model on planning problems. Although the PDDL interface is easy to use with the current task planner, it is difficult for end-users to add new tasks. In this work, we instead explore how to bridge the gap between human language and formal planning language task descriptions, reducing the need for engineer-handcrafted planning domains.
Liang et al. [5] use an LLM for generating code policies that can be executed on the robot. However, they do not account for the robot’s kinematics or geometry in the planning process, which is critical in real-world applications. Conversely, SayCan [6] and Text2Motion [7] are notable for considering geometric feasibility when planning the action sequence. PIGINET [8] instead introduces a transformer-based plan feasibility predictor to be integrated in a TAMP planner. The method fuses 6 cameras, making it harder to deploy in domestic settings. Our research diverges from these models by a multimodal structure that processes user input, visual and geometric information to output a sequence of actions that directly interfaces with the motion APIs.
Recently, applying vision-language models (VLMs) [9, 10, 11, 12] in robotics emerged as another promising direction. Robotics transformer (RT)-2 [13] employs vision to enable a more generalizable approach to its former version RT-1 [14]. Although these models report robust embodied intelligence, collecting large-scale robot-specific data is required, making it less applicable and practical. Our work also develops a data-based solution, but from a robot-agnostic perspective using text-only data.
II-B Human-Robot Interaction for Autonomy
Recent research indicates that LLMs may not fully replicate human reasoning capabilities [15]. Consequently, integrating HITL is essential for maintaining the reliability and safety of robotic systems [16]. Vemprala et al. [17] implement ChatGPT with HITL, emphasizing the dual benefits of enhanced model training through human feedback and increased safety during operations. Ren et al. [18] focus on human-robot interaction, where the robots decide whether or not to ask for human intervention. The framework, however, remains a sole task planner that evaluates task reasoning without considering the actual robot constraints. The decision to ask the human or not depends on the planner’s confidence in generating a logic solution rather than an explicit failure. Moreover, the developed algorithm also relies on a scene description from the prompt and does not have any perception input. Our approach underlines the importance of human oversight in critical decision-making problems, and facilitates a clear and intuitive interface where the human is informed during robot operation.
II-C Replanning for Failure Recovery
Failure recovery has been widely addressed in the literature for robotics planners. When it comes to LLM-based solutions, researchers typically address replanning from a second-level action check when the plan is generated [19, 20]. Vision-language models are also implemented to describe the error from the scene [21]. REFLECT [22] incorporates visual and audio sensory data to explain failures from the failed actions. This showcases the success of failure reasoning with a multimodal structure, however does not consider robotic constraints and remains a pure task planner. ReplanVLM [23] uses three different VLM agents for planning, replanning with internal and external errors, indicating that robust solution is hard to obtain by prompting one single LLM planner. CAPE [24] instead proposes re-prompting to replan upon failure with preconditions similar to classical planning methods (e.g. PDDL, STRIPS). Considering the context of task and motion planning, this method works in the form of hierarchical layers and replans upon execution failure. In addition to recovering from execution, we designed a multimodal LLM-based planner that incorporates action feasibility during generating plans, which helps to reduce the time taken for re-querying the planner.
II-D Data-efficiency in LLM-based robotics planner
LLM-based models are usually data-intensive and require substantial training datasets and computing resources to generate robust solutions. In robotics, this demand poses a significant challenge, as collecting large, diverse, and representative datasets from physical robots is often prohibitively time-consuming, labor-intensive, and costly. For instance, RT series [14, 13] and $\pi_{0}$ [12] introduces an end-to-end approach to applying transformers to robotics, requiring a large amount of data collected in months to map vision and language to actions. Similarly, Lynch et al. [25] proposes an interactive method for language-conditioned robot skills, compiling a dataset through 2.7k hours of collection on real robots. Prompt engineering, on the other hand, has shown promise in applying LLMs to robotic tasks in a zero-shot manner [26, 27, 28]. However, this is often not robust for real-world execution without considering and understanding the operating environment. Parameter-efficient fine-tuning (PEFT) [29] techniques have become a promising approach to minimize fine-tuning costs while enabling LLMs to acquire task-specific parameters. Our work employs this method to incorporate robotics knowledge to LLMs while introducing a widely applicable LLM-based framework that is easy to deploy on different robot platforms, as it does not require robot-specific data.
III Methodology
Figure 2: System overview. Our multimodal planner integrates user input, visual perception, and action feasibility score as inputs to a fine-tuned lightweight LLM. The LLM will then generate an action sequence for a real robot to perform mobile manipulation tasks.
III-A Problem Statement
[…]
The robot obtains environmental information from a vision module, a state-of-the-art object detection model, processing input from the onboard RGB-D camera. The visual detection is represented as $O$. The motion verification module assesses the feasibility score of the action during planning. In this setting, we define InteLiPlan as a real-time onboard LLM-based planner, with the ability to interact with humans in natural language and replan upon guidance. We evaluate InteLiPlan on domestic tasks, aiming at a human-centric and interpretable framework.
III-B Interactive Planner with Multimodal Perception
Our system processes multimodal inputs from the user $T_{user}$, visual observations $O$, and motion feasibility scores $f_{score}$. At the core of our system is a fine-tuned LLM that serves as the interactive planner, as illustrated in Fig. 2. Algorithm in Fig. 2 explains the internal interaction between the modules. When the user sends a command to the central planner, the request is first sent to the vision module. The vision module runs as a service that detects objects in real-time, and continuously update a list of detected objects with their corresponding positions in the vision memory. When a task is sent to the robot, the system checks if the mentioned items are found in the current observation. If the object is found, the motion verification module gets the object position and checks if the target object is reachable. For example, in our implementation, we use the reachability graph designed in R-LGP[30], which is used as a feasilibity check service for logic-geometric programming (LGP) framework [31]. This reachability graph takes in the desired end-effector position (e.g. object location $x_{obj}$) and current end-effector position and determines whether or not there is a whole-body collision-free trajectory between the two end points, regarded as $f_{score}(x_{obj},G_{map})$. In InteLiPlan, we incorporate a binary score for feasibility check, which applies, but not limited to the reachability graph in R-LGP. Lastly, we keep track of the conversation by inserting a history $T_{hist}$ to each planner call. $T_{hist}$ is initialized empty and records the initial request with the most recent failure detected by the planner.
Bringing together multimodal observations and interaction history, we formally define InteLiPlan as:
$A=\phi(T_{hist},T_{user},f_{seen}(l_{obj}),f_{score}(x_{obj},G_{map})),$ (1)
[…]
Our framework features a centralized LLM-based planner to utilize the integration and coordination between submodules. We provide a plug-and-play modular design to leverage state-of-the-art models across various research domains (e.g., object recognition) in a zero-shot way.
III-C Failure Recovery
The ability to replan task and motion upon a failure is a critical feature towards real-world applications. This can be accomplished by active exploration, or with guidance from humans. A recent paper from [32] solves a self-recovery problem with an LLM-based planner. Our work instead addresses cases where human intervention is required. For example, when the task is not clear to the robot; when the user has specific preference towards choices; or when the robot is completely blocked by its partial observation of the environment. We blend our method with the human-in-the-loop mechanism, taking advantage of human input for corner cases. By integrating human insights, InteLiPlan can navigate complex environments more effectively and perform tasks with greater precision, thereby increasing the system’s overall autonomy and operational safety.
We consider three main categories of failures: planning failure, task confusion and execution failure, as outlined in Table I. The multimodal structure allows the planner to reason around failures through feedback from the vision and motion verification modules, extracting the missing information to complete the task. When a failure occurs, our system can report to the human with a contextual description, interpreting the current observation and failure that occurred. By fine-tuning the LLM, the model can intelligently process human inputs and replan based on human guidance. For instance, vision failures due to partial observation or adverse lighting conditions can be mitigated with human insights, leveraging the planner’s capability to maintain a record of initial commands through the internal dialogue $T_{hist}$ and to replan upon system failure as necessary. This capability significantly enhances transparency, robustness, and reliability. By injecting fine-tuning data, the output is guaranteed to consider necessary observations before executing actions.
TABLE I: Cases that require human interaction.
Category Case Examples
Planning failure Vision failure Failed to find object
Feasibility failure Unable to reach object
Task confusion Ambiguous reference Found more than 1 specified object
Ambiguous task The user asks for ”something to drink”
Execution failure Action failure Ask for recovery
III-D Text-only Fine-Tuning Data for Robotics Planner
Recently, prompt engineering has become popular for implementing ready-to-use LLM-based planners, but poses a risk of plan reliability when executing on embodied agents. In addition, outputs from such planners often need translation into robotic control commands. Our method implements few-shot fine-tuning to inject the multimodal structure to the LLM, as well as inform the planner of the desired output format that enables direct calls to the action API. This approach also lays out the context and reduces the need for additional conversion, which uses computing resources that can prevent real-time planning.
Figure 3: Examples of the multimodal LLM-based planner. (a) presents a no-failure case for the given command. (b) and (c) depict the failure reasoning ability of the model considering the inputs, for vision and feasibility failures respectively. (d) showcases the ability to recover from the failure in (b) with human instruction.
To fine-tune the LLM model, we collected a pure-text customized dataset that includes scenarios both with and without failure recovery, where the interactions in failure cases are limited to a maximum of two rounds. The dataset is formulated as $D=\{T_{hist},T_{user},f_{seen},f_{score},A\}$. We loop through a list of objects and commands to generate the command and expected sequence of actions. Visual observation $f_{seen}$ and feasibility score $f_{score}$ will be processed by the fixed visual module and feasibility module to generate textual signals. This text-based-only fine-tuning approach enables the model to easily adapt to different robots without further modifications or re-parameterization.
Our methodology employs Low-Rank Adaptation (LoRA) [33] to fine-tune LLM efficiently. Instead of updating all model parameters, LoRA introduces trainable low-rank matrices into key projection layers, significantly reducing computational and memory overhead. LoRA is one of the PEFT [29] techniques that has also been proven to be superior over full fine-tuning as a parameter-efficient fine-tuning method in [34]. We implement supervised fine-tuning with the collected dataset, where $\{T_{hist},T_{user},f_{seen},f_{score}\}$ is the input, and $A$ is the desired output, with cross-entropy loss. This approach enables efficient adaptation of the model to domain-specific tasks while maintaining scalability.
IV Experimental Results
IV-A Evaluation of Robustness Across Different Failures
In the first set of experiment, we evaluate our approach using the Toyota Human Support Robot (HSR) in a ‘fetch me’ task. The environmental setup includes 10 seen objects and 20 unseen objects, and the planner is evaluated on 4 seen request commands, 23 unseen request commands, and replan with 9 seen guidance commands, and 18 unseen guidance commands. The unseen commands represent the variants in human-like conversation. The vision module receives inputs from the RGB-D camera of the HSR, then processes object recognition with YOLO [35] and uses pose estimation to obtain object locations. For feasibility verification, we use the reachability graph from [30], a sampling-based approach that efficiently checks for paths from a starting point to a goal while ensuring feasible configurations for the robot reaching towards the target object. To verify applicability, the LLM model for our approach is fine-tuned from Mistral 7B [36] and Deepseek 8B [37], state-of-the-art lightweight LLM models, which is suitable for onboard deployment. The fine-tuning data includes 300 samples covering all failure categories outlined in Table I. Before the execution, the robot is provided with a map of the room, e.g. large furniture like tables or drawers, and a list of motion APIs that are shared between the planner and the controller.
To assess the performance of the proposed method in this experiment, we use three metrics. 1) Task planning success rate: the ratio of the no-failure trials where the robot successfully generates the action sequence by understanding the assigned task. 2) Failure explanation success rate: the ratio of the trials where the robot successfully generates the reason for not being able to find the action sequence. 3) Failure recovery success rate: the ratio of the trials where the robot successfully generates the recovery plan in response to human guidance.
We compare Mistral-Ours and Deepseek-Ours with few-shot prompting on Mistral - Mistral-Prompt and Deepseek - Deepseek-Prompt. The samples are in the same form as the dataset for fine-tuning, as discussed in Section III-D.
IV-A1 Comparison of Different Methods in Feasible Tasks
In our first experiment, we assume successful outcomes in both object detection and robotic motion feasibility. Here, the planner is solely required to generate the appropriate action sequences based on human inputs. An example of such conversations is provided in Fig. 3a.
Despite explicit instructions to produce solely an action sequence, Mistral-Prompt and Deepseek-Prompt often generate extraneous text. For a fairer comparison, we only extracted the outputted action sequence when determining the success rate, and ignored the lengthy explanation and slightly misused words (e.g., ‘pick up’ vs ‘pick’) for the action API interface. As presented in the w/o failure column of Table II, the results demonstrate a superior performance of our models, achieving 100% success in scenarios with seen commands and objects, and maintaining high effectiveness even with unseen objects or commands. In contrast, Mistral-Prompt and Deepseek-Prompt experienced performance drops in scenarios involving unseen elements, with more sensitivity over unseen textual commands. It is also observed that Deepseek-Prompt behaves better than Mistral-Prompt, while Deepseek-Ours has slightly lower success rates than Mistral-Ours.
IV-A2 Evaluation of Robustness in Failure Explanation
In this test, we break down the evaluation of failure explanation in the failure recovery process, investigating the ability of the model to explicitly describe what hinders the process of planning actions. The failure experiments include all categories outlined in Table I, with examples of ‘planning failure’ detection in Fig. 3b,c. We only include ‘execution failure’ in failure recovery experiments (Section IV-A3), as it is a failure description that is triggered from the controller level.
Figure 4 shows that our method achieved near-perfect success rates in scenarios with seen commands and seen/unseen objects, underscoring its reliability in familiar settings. It is observed that ‘Task confusion’ failure detection obtains lower success rates than multimodal failure. This is mainly due to the ‘ambiguous task’ experiments, where the robot fails to distinguish object categories (e.g. fruits, drinks) from the list of detected objects. Our method increases the success rates of failure explanation by more than twice in these cases, proving that fine-tuning triggers the domain-specific understanding of the robots for which factor (e.g. user preference) to consider during planning.
IV-A3 Evaluation of Robustness in Failure Recovery
This section evaluates the ability of our framework’s capacity to effectively replan from human instructions. For each trial, we assume a detected failure in <history>, and expect the models to output corresponding action to the user inputs.
Fig. 5 demonstrates our framework’s performance in replanning in various scenarios. While ‘task confusion’ mainly processes user preference (e.g. user prefers the robot to get a coke to a 7up), ‘planning failure’ and ‘execution failure’ cases assess the ability of mapping language to the list of pre-defined motions (e.g. output ‘open cupboard’ in response to ‘the object is in the cupboard’). The lower success rates in recovering from ‘execution failure’ are due to the longer sequence of recovery than what was provided in the fine-tuning data. In general, prompting methods fails to embody LLM with sufficient robot-domain understanding. Our method significantly increases the performance of LLM models in robotic task failure recovery, indicating robustness in real-world conditions where adaptability and responsiveness to failures are crucial for operational success.
The w/o failure column of Table II reports success rates in scenarios where the planner must succeed in the two-step process of failure detection followed by failure recovery. The results reveal a significantly increased gap between prompt-based models and our approach as opposed to w/o failure cases, underscoring the effectiveness of our method in handling multi-step interaction and ensuring task completion.
TABLE II: Success Rates (%) of Task Planning.
Scenario Method w/o failure w/ failure
Seen cmd + Seen obj Deepseek-Prompt 100 100
Deepseek-Ours 100 100
Mistral-Prompt 100 52
Mistral-Ours 100 100
Seen cmd + Unseen obj Deepseek-Prompt 90 35
Deepseek-Ours 100 86
Mistral-Prompt 90 29
Mistral-Ours 100 90
Unseen cmd + Seen obj Deepseek-Prompt 98.95 54
Deepseek-Ours 100 93
Mistral-Prompt 86.32 49
Mistral-Ours 100 94
Unseen cmd + Unseen obj Deepseek-Prompt 95.79 39
Deepseek-Ours 97.89 93
Mistral-Prompt 78.95 36
Mistral-Ours 100 95
Figure 4: Success rates of failure explanation.
Figure 5: Success rates of failure recovery.
IV-B Evaluation of Scalability Across Tasks
In this experiment, we evaluate the scalability of the approach by 101 task instructions from SayCan [6]. To accommodate the task variants, we fine-tuned Mistral with the tasks outlined in Table III, which covers the expected actions required for the dataset solutions.
This set of experiments added a metric of Execution success rate, which is the ratio of the trials where the robot successfully executes the planned actions. Given that SayCan [6] does not have failure recovery ability, we only report the result of cases without failure recovery. Generative result from PaLM 8B [38] is also reported as a lightweight LLM baseline.
TABLE III: Tasks with Corresponding Expected Outcome.
Case Expected plan
Pick object Pick object
Go to destination Go to destination
Fetch me Pick object, Go back, Place object
Put away Pick object, Go to destination, Place object
Put in drawer Open drawer, Pick object, Place in drawer, Close drawer
Table IV presents the tested result of the frameworks with the dataset from Saycan. As a 540B model, the SayCan result explains its excellent decision-making capability with throughout understanding of the world. Its trained affordance value provides sufficient embodied knowledge, ensuring high success rate in execution. Besides, we note that prompting with PaLM 8B only successfully plans 38% cases. Mistral-Prompt with our modular structure helps the LLM model to gain embodied intelligence, with the planning success rate to 59%. It is observed that Mistral-Prompt’s failure cases come from its lack of sense of the operating environment. For example, some of the results tell the robot to ‘go to store’ to pick something up, despite being told that it is working in a domestic environment. It is proven that with 7B parameters, our InteLiPlan obtains similar results in comparison to the state-of-the-art 504B SayCan model with 83% planning success rate.
The execution success rate remains relative to the planning success rate in Mistral-Prompt and Mistral-Ours, given the same vision and reachability modules check for the system. Having the reachability graph checking whole-body feasibility explicitly instead of using skill probabilities like SayCan, our approach also increases the chances of successfully executing the plan once generated. This is shown by the reduced difference in plan and execute success rates between SayCan and our pipeline. On the other hand, as we use a sampling-based approach, there is a risk of not finding a path even if there is one exists, leading to failures from the planning level. This is where the failure recovery comes in. Another solution is to use more samples in the graph such that they sufficiently provide the reachability knowledge. Since InteLiPlan is a plug-and-play system, it can seamlessly integrate future state-of-the-art feasibility modules.
TABLE IV: Plan and Execution Success Rates (%) of the methods for 101 SayCan task descriptions.
Methods Plan Execute
PaLM-8B 38 n/a
PaLm-504B SayCan 84 74
Mistral-7B Prompt 59 58
Mistral-7B Ours 83 82
IV-C Evaluation of Cross-embodiedment with Real-robot Experiments
We validate the cross-embodiment feature of the system on two different physical robot platforms that are different in both camera setups and the number of degrees of freedom (DoF). In both robot setups, InteLiPlan is processed entirely onboard with a single NVIDIA Geforce RTX 2070, verifying onboard computing capability. We tested a pre-trained InteLiPlan model on both platforms:
  • Toyota Human Support Robot (HSR): 11-DoF robot. The vision module takes the signal from the Asus Xtion PRO Live stereo camera on the head of the robot.
  • Anymal D-Unitree Z1 (Anytree): 12-DoF robot (where the Anymal D is simplified with a planar base). The vision module takes the signal from the Realsense D435 on the wrist of the end effector.
We implemented 3 sets of experiments, with 10 trials each:
  • No-failure task: the robot is asked to complete a task by processing a high-level command. In this task, we guarantee that the object is within the observation and is reachable.
  • 1-step failure task: we implement the planning failure from vision, where the targeted object is not within its observation, the robot is expected to report the failure to the user and replan upon guidance.
  • Multi-step failure task: the robot must detect failures and replan twice to be able to complete the task. Specifically, we implement the failure from vision first, then after the targeted object is found, it is out of reach and the robot should ask again to complete the task.
TABLE V: Success Rates (%) of Task Planning.
Robot w/o failure 1-step failure Multi-step failure
HSR 100 100 90
Anytree 100 100 90
Table V recorded the success rates of InteLiPlan in generating the correct sequences of action upon the corresponding case. The success rates were also affected by the success of vision and feasibility modules, leading to the failed cases. The result demonstrates that with one model, planned action sequences guarantee executability across embodiments. Fig. 1 and Fig. 6 showcase InteLiPlan results in real-world settings, which include tasks with and without failure recovery. The multi-step failure verifies that our method enables autonomy with seamless end-user intervention.
((a)) Failure from feasibility
((b)) Failure from vision
Figure 6: Demonstration of the interactive failure recovery capability on the physical HSR.
Table VI records the averaged planning time of the system over the trials, which is divided into vision, feasibility and planner queries processing time. In total, the model can plan an executable action sequence in less than 7s with onboard resources. The feasibility check module, which takes up the longest processing time, can be relaxed with fewer nodes in the reachability graph. Notably, using YOLO instead of VLM for vision input significantly contributes to real-time and onboard processing capability of the pipeline.
TABLE VI: Breakdown of average onboard planning time(s).
Robot Vision query Feasibility query Planner query
HSR 6e-6 5 1.5
Anytree 8e-6 5.5 1.3
The experiment validates the cross-embodiedment feature that requires no extra finetuning, further strengthening that InteLiPlan’s planning capability is universal to different robots, and also different camera views, as opposed to vision-language-action models [39].
V Conclusion
We presented InteLiPlan, an interactive lightweight LLM-based robotics planner for reliable and robust autonomy in domestic environments. Our framework employs a conversation-like format between internal modules and facilitates human-robot interaction, allowing the system to reason about failures through a multimodal input formulation. By incorporating a human in the loop, with human-like text understanding capabilities of LLMs, the robot gains the ability to replan based on instructions, effectively utilizing human input for dynamic adjustments. We performed extensive evaluations to investigate common failures that happen in real scenarios, where human instruction is required. The interactive behaviors guarantee interpretability and reliability for robots operating in domestic environments.
This paper reduces the effort of defining traditional task planning domains in robotics, provides an intuitive, robot-independent data structure and boosts applicability in resource-constrainted applications. Notably, deploying our approach with the lightweight Mistral 7B model achieves both comparable results with the SOTA baseline and –notably to our approach– real-time onboard computing. Our system was validated on both the physical HSR and Anymal D-Unitree Z1 robots, verifying the cross-embodiedment capability across different DoFs and camera views.
Limitations and Future Work - Our framework is structured as a multimodal system with a feasibility check, however, it does not have a motion-level replanning feature. Incorporating low-level reactions together with high-level replanning capabilities in the form of a dual process can potentially improve failure recovery speed. Furthermore, the sampled path formed by the reachability graph can be further developed into a motion planning subpart of our approach, resulting into a full task and motion planning stack.